A company owes a bank 2 payments. The first is for $ 600,000 and expires in 3 months, while the second is for $ 750,000 and expires in 5 months. If the company wants to pay off the debt by means of a one-time payment of $ 1,346,981.50, on what date should the payment be made, if the agreed interest rate is 27% per year / biweekly?
Answer #1

rate compounded monthly=((1+27%/26)^(26/12)-1)*12=27.16%

600000/(1+27.16%/12)^3+750000/(1+27.16%/12)^5=1346981.50/(1+27.16%/12)^n

=>n=4.00

So, payment must be made in 4 months